Solve the eqaution by the squared root property.
(4x+4) squared = 64

Answer by checkley77(12844)

(4x+4)^2=64
4x+4=8
4x=8-4
4x=4
x=1 answer.
proof
(4*1+4)^2=64
(4+4)^2=64
8^2=64
64=64